A linear model was fit to predict weekly Sales of frozen pizza (in pounds) from the average Price ($/unit) charged by a sample of stores in a city in 39 weeks over a three-year period. The average Sales were 61,802 pounds (SD=10,070 pounds), and the correlation between Price and Sales was −0.535. If the Price in a particular week was two SD higher than the mean Price, how much pizza would you predict was sold that week?
Answer #1

cause Price in a particular week was two SD higher than the mean Price ; therefore zx =2

here z score for sales zy =r*zx   =-0.535*2=-1.07

hence corresponding sales for the week =mean +z*std deviation =61802-1.07*10070=$51027.1